OpenStudy (anonymous):

I think it's A because you divide on both sides, but can someone help me out? I'll fan/medal The steps below show the incomplete solution to find the value of y for the equation 6y − 3y − 5 = −3 + 23: Step 1: 6y − 3y − 5 = −3 + 23 Step 2: 6y − 3y − 5 = 20 Step 3: 3y − 5 = 20 Which of these is most likely the next step? 3y = 4 3y = 15 3y = 25 3y = 5

OpenStudy (phi):

Step 3: 3y − 5 = 20 the next step would be to "move" the 5 to the right hand side

OpenStudy (anonymous):

@phi Mhm, I did that, so I divide 20 / 5 right? Because it's a multiplication problem with 3y... or do I add?

OpenStudy (phi):

you would add +5 to both sides 3y − 5 +5 = 20+5 if you divide each side by 5 you would get this \[ \frac{3y-5}{5}= \frac{20}{5} \\ \frac{3y-5}{5}= 4\] which does not look promising...

OpenStudy (anonymous):

@phi Thank you!